### v1

@@ Abstract.
The notion of sectional-hyperbolicity is a weakened form of hyperbolicity introduced for vector fields in order to understand the dynamic behavior of certain higher-dimensional systems such as the multidimensional Lorenz attractor. In this paper we explore the questions proposed in [Math. Z., 298 (2021), 469-488] and we provide an affirmative answer by proving that $C^{1}$-generic non-trivial sectional-hyperbolic chain-recurrent class is robustly a homoclinic class. We then apply our findings to the dynamics of star flows providing a positive answer to Conjecture 1.2 in [Tran. Amer. Math. Soc. 376 (2023), 6845-6871]. We remark that our arguments do not assume Lyapunov stability.
@@ 1. Introduction
In this work, we explore this question and some interesting implications. One important aspect of all of the aforementioned results is that they assert the non-existence of aperiodic classes under good generic conditions. Indeed, the existence of aperiodic classes represents a major obstacle in several important conjectures. See for instance the weak Palis conjecture which asserts that any $C^{1}$-system is approximated by either a Morse-Smale system or a system containing a hyperbolic horseshoe. This conjecture was verified for diffeomorphisms in [12], for nonsingular flows in [32] and for singular three-dimensional flows in [16].
On the other hand, in [13] and [17] the existence of periodic orbits for generic chain-recurrence classes was obtained under the assumption that the class is Lyapunov stable. Also, in [28] a dichotomy for chain-recurrence classes of generic star flows was obtained: either they have periodic orbits and positive entropy or they are sectional-hyperbolic aperiodic classes, and hence non-Lyapunov stable, with zero entropy. This evidences that obtaining periodic orbits for generic chain-recurrence classes is challenging, without Lyapunov stability. In this work, we shall see that in many cases, sectional-hyperbolicity is enough to rule out aperiodic classes.
We now state our main results. Let us begin by recalling the concept of sectional-hyperbolic set. We say that a compact invariant set $\Lambda$ has a dominated splitting if there is a continuous invariant splitting $T_{\Lambda}M=E\oplus F$ (with respect to the tangent flow $\Phi_{t}=DX_{t}$, $t\in\mathbb{R}$) and constants $K,\lambda>0$ satisfying the relation
@@
Next, let us clarify the notion of a robustly transitive sectional-hyperbolic chain-recurrent class. It is immediate to verify that $\operatorname{Crit}(X)\subset CR(X)$, so the chain-recurrent class of any critical element is well defined. Moreover, by the previous remark and the hyperbolic lemma (see Lemma 2.1), every critical element of a sectional-hyperbolic chain-recurrent class is hyperbolic. If $\gamma_{X}$ is a critical element of $X$ and $C(\gamma_{X})$ is sectional-hyperbolic, then for every $Y\in\mathcal{X}^{1}(M)$ sufficiently $C^{1}$-close to $X$, there exists a continuation $\gamma_{Y}$ of $\gamma_{X}$. The corresponding chain-recurrent class $C(\gamma_{Y})$ is called the continuation of $C(\gamma_{X})$. We say that the chain-recurrent class $C(\gamma_{X})$ is robustly a homoclinic class if there exists an open $C^{1}$-neighborhood $\mathcal{U}$ of $X$ such that for every $Y\in\mathcal{U}$, the continuation $C(\gamma_{Y})$ is a homoclinic class. The first main result of this paper is the following:
Theorem A.
@@
There exists a $C^{1}$-generic set $\mathcal{R}\subset\mathcal{X}^{1}(M)$ such that if $\Lambda$ is a non-trivial sectional-hyperbolic chain-recurrent class, then it is robustly a homoclinic class.
Notice that in the previous result we dropped out any stability assumption on $\Lambda$. In particular, Theorem A is a direct improvement of [27, Corollary E]. Also, an implicit fact behind the statement of Theorem A is that any non-trivial sectional-hyperbolic chain-recurrent set within $\mathcal{R}$ must contain periodic orbits; in other words, sectional-hyperbolicity is enough to rule out the existence of aperiodic classes.
Another application of Theorem A concerns the setting of star flows. Recall that a vector field $X\in\mathcal{X}^{1}(M)$ is star if there exists a neighborhood $\mathcal{U}$ such that all critical elements of every $Y\in\mathcal{U}$ are hyperbolic. Star systems play a central role in stability theory. For diffeomorphisms, it was proved in [23, 21] that $\Omega$-stability implies the star condition, and later in [19] that the star condition is equivalent to Axiom A with the no-cycle condition. For regular vector fields, the same equivalence was obtained in [15].
Beyond stability, singular star flows are known to exhibit rich dynamics, which has motivated substantial research. In [33], it was conjectured that $C^{1}$-generic star flows and sectional-hyperbolicity should be intrinsically related; namely, they were expected to have finitely many chain-recurrent classes, each sectional-hyperbolic. A partial confirmation was obtained in [30] under the assumption of Lyapunov stability. However, [7, 22] disproved half of the conjecture by showing that $C^{1}$-generic star flows need not be sectional-hyperbolic, thus opening a new line of research with the introduction of multi-singular hyperbolic sets. Nevertheless, the other half of the conjecture remains open:
Conjecture 1.3 (Conjecture 1.2 in [28]).
$C^{1}$-generic singular star flows do not admit singular aperiodic classes. Consequently, they have only finitely many chain-recurrent classes, all of which are homoclinic classes of periodic orbits.
A partial answer was given in [28], proving that generic star flows admit finitely many Lyapunov chain-recurrent classes and establishing a dichotomy for the non-Lyapunov stable case, based on entropy. As a consequence of our techniques, we can fully confirm the conjecture:
Theorem B.
Let $M$ be an $n$-dimensional manifold with $n\geq 4$. There exists a $C^{1}$-generic set $\mathcal{R}\subset\mathcal{X}_{*}^{1}(M)$ such that for every $X\in\mathcal{R}$ and every non-trivial chain-recurrent class $\Lambda$ of $X$, the set $\Lambda$ has positive topological entropy, contains a periodic orbit, and is isolated. Consequently, $C^{1}$-generic singular star flows have only finitely many chain-recurrent classes, all of which are either trivial or homoclinic classes of hyperbolic periodic orbits.
@@
The remainder of this text is organized as follows: In Section 2, we present a collection of preliminary results concerning sectional-hyperbolic chain-recurrence classes. In Section 3, we prove that any chain-recurrent class associated with $X\in\mathcal{R}$ is robustly periodic, which constitutes a key step toward proving Theorem A. Finally, Section 4 is devoted to proving the main results of this article.
@@ 2. Preliminary results on Sectional-Hyperbolic chain-classes
In this section, we collect some tools that will be used in our analysis. Let us begin by recalling that an important fact about chain-recurrence classes is that they are chain transitive: If $C$ is a chain-recurrent class, then for any pair of points $x,y\in C$ the point $y$ is chain attainable from $x$ through points in $C$, i.e., for every $\varepsilon,T>0$, there is an $(\varepsilon,T)$-chain $(x_{i},t_{i})_{i=0}^{n}$ such that $x_{0}=x,x_{n}=y$ and $x_{i}\in C$ for any $i=0,\ldots,n$ (see for instance [14, Proposition 2.1]).
@@
Notice that by Lemma 2.2 and Remark 2.3, we can consider a useful tool for proving Theorem A, coming from [28]. More precisely, let $E^{cs}\oplus E^{u}=E^{s}_{\sigma}\oplus F_{\sigma}$, where $E^{cs}=E_{\sigma}\oplus G$ and $F_{\sigma}=G\oplus E^{u}_{\sigma}$, with $G\subset F$ and $\text{dim }G=1$, be the hyperbolic splitting of $T_{\sigma}M$. Up to a change of coordinates, we think $\sigma$ as the origin in $\mathbb{R}^{n}$, and the subspaces $E^{*}$, =cs,u*=cs,u, are perpendicular. As in the aforementioned reference, for a small neighborhood $U_{\sigma}$ of $\sigma$ and $x\in U_{\sigma}$, we denote by $x^{cs}$ and $x^{u}$ its distances to $E^{u}$ and $E^{cs}$ respectively. In this way, for $\alpha,\beta>0$ define the $\alpha$-cone on the manifold $D_{\alpha}^{cs}(\sigma)$ as
$D_{\alpha}^{cs}(\sigma)=\{x\in U:x^{u}<\alpha x^{cs}\}.$
$N_{x}=\{v\in T_{x}M:\langle v,X(x)\rangle=0\}.$
$\psi^{*}_{t}(v)=\frac{\|X(x)\|}{\|X(X_{t}(x))\|}\psi_{t}(v).$
In [6] it is showed that a partially hyperbolic splitting $T_{\Lambda}M=E\oplus F$ of a compact invariant set $\Lambda$ induces a partially hyperbolic splitting $T_{x}N_{x}=\mathcal{E}_{x}\oplus\mathcal{F}_{x}$ of the normal subspace $N_{x}$ for any $x\in\Lambda\setminus Sing(X)$, where $\mathcal{E}_{x}$ and $\mathcal{F}_{x}=\pi_{x}(E_{x})$ for every $x\in\Lambda$. Indeed, since $F_{x}$ contains the flow direction, it follows that $\pi(v)\in F_{x}$ for any $v\neq X(x)$ in $F_{x}$, hence $\mathcal{F}_{x}\subset F_{x}$ because $\pi$ is linear. So, since the splitting $E\oplus F$ is dominated for $\Phi_{t}$, we deduce the dominance of the splitting $\mathcal{E}\oplus\mathcal{F}$. The uniform contracting property for $\pi(E)$ follows because $\|\pi(v)\|\leq\|v\|$ for every $v\in E$. It is important to observe that the above property remains valid for the rescaled linear Poincaré flow $\psi^{*}_{t}$.
Recall the notion of $(\eta,T,E)$-contracting points.
Definition 2.4.
Let $X\in X^{1}(M)$, $\Lambda$ a compact invariant set of $X$, and $\mathcal{E}\subset N_{\Lambda\setminus Sing(X)}$ an invariant bundle of the rescaled linear Poincaré flow $\psi^{*}_{t}$. For $0<\eta<1$ and $T_{0}>0$, we say that a point $x\in\Lambda\setminus Sing(X)$ is $(\eta,T_{0})$-contracting for the bundle $\mathcal{E}$ if there is a partition $0=t_{0}<t_{1}<t_{2}<\cdots<t_{k}<\cdots$, where $t_{i+1}-t_{i}\in[T_{0},2T_{0}]$ for every $i=0,1,\ldots$ such that
$\prod_{i=0}^{k-1}\|\psi^{*}_{t_{i+1}-t_{i}}|_{\mathcal{E}(X_{t_{i}}(x))}\|\leq\eta^{k},\quad\forall k\geq 1.$
Similarly, $x\in\Lambda\setminus Sing(X)$ is said to be $(\eta,T_{0})$-expanding for the bundle $\mathcal{F}$ if it is $(\eta,T_{0})$-contracting for X-X.
We are now ready to state the result mentioned above.
Lemma 2.5 (Lemma 3.11 in [28]).
Let $\sigma$ be a Lorenz-like singularity. Then for $r>0$ small enough, for every $\lambda\in(0,1)$ and $T_{0}>0$ there exists $\alpha_{2}>0$ such that for all $\alpha<\alpha_{2}$, if $y\in B_{r}(\sigma)\cap\overline{D_{\alpha}^{cs}(\sigma)}$ is a $(\lambda,T_{0})$-expanding time on its orbit, then $W_{loc}^{u}(y)\pitchfork W_{loc}^{cs}(\sigma)\neq\emptyset$.
Next, we recall a remarkable theorem due to Liao regarding the existence of stable (unstable) manifolds for contracting (expanding) points.
Theorem 2.6.
Let $X\in X^{1}(M)$, $\Lambda$ a compact invariant set of $X$. Given $\eta\in(0,1)$, $T_{0}>0$, there is $\delta=\delta(\eta,T_{0})>0$ such that if $x$ is $(\eta,T)$-contracting for the bundle $E\subset N$, then there is a $C^{1}$ map $\kappa:E_{x}(\delta\|X(x)\|)\to N_{x}$ such that
  • $d_{C^{1}}(\kappa,id)<\varepsilon$.
  • $\kappa(0)=0$.
  • $W_{\delta\|X(x)\|}^{s}(x)\subset W^{s}(\mathcal{O}(x))$, where $W_{\delta\|X(x)\|}^{s}(x)=exp_{x}(\text{Image}(\kappa))$.
A similar result holds for $(\eta,T_{0})$-expanding points for the bundle $F\subset N$.
Finally, we end this subsection by showing that sectional-hyperbolicity of a compact invariant set can be extended to the maximal invariant set of a neighborhood of it.
@@ 3. Robustly Periodic chain-recurrence classes
In this section, we shall derive one of the main results of this paper. Namely, we prove that $C^{1}$-generically every sectional-hyperbolic non-trivial chain-recurrent class robustly periodic. The statement is as follows:
@@
There exists a $C^{1}$-generic set $\mathcal{R}\subset\mathcal{X}^{1}(M)$ such that if $\Lambda$ is a non-trivial sectional-hyperbolic chain-recurrent class then it is robustly periodic, that is, there is an open neighborhood $\mathcal{U}$ of $X$ such that for every $Y\in\mathcal{U}$, one has $Per(Y|_{\Lambda_{Y}})\neq\emptyset$, where $\Lambda_{Y}$ denotes the continuation of the chain recurrent class $\Lambda$ for $X$.
In order to prove Theorem 3.1, we shall show first that $C^{1}$-generically every sectional-hyperbolic non-trivial chain-recurrent class actually contains a periodic orbit, and therefore it is a homoclinic class. As a direct consequence, it has positive topological entropy.
@@
Next, consider the following generic set that is obtained by combining some items from [13, Proposition 2.1 ] and [17, Lemma 3.12].

### v2

@@ Abstract.
The notion of sectional-hyperbolicity is a weakened form of hyperbolicity introduced for vector fields in order to understand the dynamical behavior of certain higher-dimensional systems such as the multidimensional Lorenz attractor. In this paper we address the questions proposed in [Math. Z., 298 (2021), 469-488] and we provide a partial answer by proving that a $C^{1}$-generic non-trivial sectional-hyperbolic chain-recurrent class, not necessarily Lyapunov stable, satisfies a trichotomy: it is either a homoclinic loop, a union of saddle connections between singularities, or it is robustly a homoclinic class.
@@ 1. Introduction
In this work, we explore this question, and we give a partial answer to this question. In order to state our main results, let us begin by recalling the concept of sectional-hyperbolic set. We say that a compact invariant set $\Lambda$ has a dominated splitting if there is a continuous invariant splitting $T_{\Lambda}M=E\oplus F$ (with respect to the tangent flow $\Phi_{t}=DX_{t}$, $t\in\mathbb{R}$) and constants $K,\lambda>0$ satisfying the relation
@@
Next, let us clarify the notion of robustly homoclinic class. it is immediate to verify that $\operatorname{Crit}(X)\subset CR(X)$, so the chain-recurrent class of any critical element is well defined. Moreover, by the previous remark and the hyperbolic lemma (see Lemma 2.1), every critical element of a sectional-hyperbolic chain-recurrent class is hyperbolic. If $\gamma_{X}$ is a critical element of $X$ and $C(\gamma_{X})$ is sectional-hyperbolic, then for every $Y\in\mathcal{X}^{1}(M)$ sufficiently $C^{1}$-close to $X$, there exists a continuation $\gamma_{Y}$ of $\gamma_{X}$. The corresponding chain-recurrent class $C(\gamma_{Y})$ is called the continuation of $C(\gamma_{X})$. We say that the chain-recurrent class $C(\gamma_{X})$ is robustly a homoclinic class if there exists an open $C^{1}$-neighborhood $\mathcal{U}$ of $X$ such that for every $Y\in\mathcal{U}$, the continuation $C(\gamma_{Y})$ is a homoclinic class. The main result of this paper is the following:
Main Theorem.
@@
There exists a $C^{1}$-generic set $\mathcal{R}\subset\mathcal{X}^{1}(M)$ such that any non-trivial sectional-hyperbolic chain-recurrent class $\Lambda$ satisfies the following trichotomy:
  1. (1)
    $\Lambda$ is a homoclinic loop.
  2. (2)
    $\Lambda$ consists of saddle connections between singularities of $\Lambda$.
  3. (3)
    $\Lambda$ is robustly a homoclinic class.
Notice that in the main result above we are not assuming the class is Lyapunov stable. Indeed, if the class is Lyapunov stable, then [14] implies that the trichotomy is reduced to case (3) of the Main Theorem. Removing this assumption imposes substantial obstacles to the analysis of generic chain-recurrent classes (see [28, 18, 29]). In this direction, our result allows one to recover robust homoclinic classes, provided the original classes are neither homoclinic loops nor a saddle connections. This shows that sectional-hyperbolicity, rather than Lyapunov stability, is the fundamental mechanism for the emergence of homoclinic structure in the generic setting.
@@
The remainder of this text is organized as follows: In Section 2, we present a collection of preliminary results concerning sectional-hyperbolic chain-recurrence classes. In Section 3, we prove that any chain-recurrent class associated with $X\in\mathcal{R}$ which is not a homoclinic loop or a saddle connection is robustly periodic, which constitutes a key step toward proving the Main Theorem. Finally, Section 4 is devoted to proving our main result.
@@ 2. Preliminary results on Sectional-Hyperbolic chain-classes
In this section, we collect some tools that will be used in our analysis. Let us begin by recalling that an important fact about chain-recurrence classes is that they are chain transitive: If $C$ is a chain-recurrent class, then for any pair of points $x,y\in C$ the point $y$ is chain attainable from $x$ through points in $C$, i.e., for every $\varepsilon,T>0$, there is an $(\varepsilon,T)$-chain $(x_{i},t_{i})_{i=0}^{n}$ such that $x_{0}=x,x_{n}=y$ and $x_{i}\in C$ for any $i=0,\ldots,n$ (see for instance [15, Proposition 2.1]).
@@
We end this subsection by showing that sectional-hyperbolicity of a compact invariant set can be extended to the maximal invariant set of a neighborhood of it.
@@ 3. Robustly Periodic chain-recurrence classes
In this section, we establish one of the main results of the paper. More precisely, we prove that for $C^{1}$-generic vector fields, every non-trivial sectional-hyperbolic chain recurrent class which is not reduced to either a homoclinic loop or a saddle connection between singularities is robustly periodic. This result is stated as Theorem 3.1 below and constitutes a key ingredient in the proof of Theorem A, as it allows us to rule out aperiodic behavior for this chain classes in the generic setting.
@@
There exists a $C^{1}$-generic set $\mathcal{R}\subset\mathcal{X}^{1}(M)$ such that if $\Lambda$ is a non-trivial sectional-hyperbolic chain-recurrent class which is not reduced to either a homoclinic loop or a saddle connection between singularities, then it is robustly periodic, that is, there is an open neighborhood $\mathcal{U}$ of $X$ such that for every $Y\in\mathcal{U}$, one has $Per(Y|_{\Lambda_{Y}})\neq\emptyset$, where $\Lambda_{Y}$ denotes the continuation of the chain recurrent class $\Lambda$ for $X$. In particular, $\Lambda$ is a homoclinic class.
In order to prove Theorem 3.1, we shall show first that $C^{1}$-generically every sectional-hyperbolic non-trivial chain-recurrent class which is not reduced to either a homoclinic loop or a saddle connection between singularities actually contains a periodic orbit, and therefore it is a homoclinic class. As a direct consequence, it has positive topological entropy.
@@
Next, consider the following generic set that is obtained by combining some items from [14, Proposition 2.1 ] and [18, Lemma 3.12].
@@
Remark 3.3.
Next, we describe the homoclinic structure of $\Lambda$ when it does not contain singularities.
Lemma 3.4.
Let $\mathcal{R}$ be the residual set of Lemma 3.2, and let $\Lambda$ be a sectional-hyperbolic chain recurrent class without singularities. Then, it is a homoclinic class.
Proof.
First, let $U$ be the neighborhood given by Lemma 2.4 and let
$\Lambda^{\prime}=\Lambda\cup\bigcup_{n\geq 1}\gamma_{n}.$
Claim 1.
$\Lambda^{\prime}$ is a compact and invariant subset of $U$.
Proof.
First, observe that $\Lambda^{\prime}\subset U$ and $\Lambda^{\prime}$ is invariant by definition. Next, let $\{x_{k}\}_{k\geq 1}\subset\Lambda^{\prime}$ be a sequence such that $x_{k}\to x\in M$, when $k\to+\infty$. If $x_{k}\in\Lambda$ for sufficiently large $k$, then $x\in\Lambda$, because $\Lambda$ is compact. Similarly, if $x_{k}\subset\gamma_{n}$ for sufficiently large $k$ and some $n\geq 1$ fixed, then $x\in\gamma_{n}\subset\Lambda^{\prime}$. It remains to verify the case where there are sequences $k_{l},n_{l}\to\infty$, $n_{l}\neq n_{l^{\prime}}$, for $l\neq l^{\prime}$ and $x_{k_{l}}\subset\gamma_{n_{l}}$. In this case, since $d^{H}(\gamma_{n},\Lambda)\to 0,$ we have $d(x_{k_{l}},\Lambda)\to 0$. Hence, $x\in\Lambda\subset\Lambda^{\prime}$ and the claim holds. ∎
Since $\Lambda^{\prime}\subset U$, Claim 1 and the hyperbolic lemma imply that $\Lambda^{\prime}$ is a compact invariant hyperbolic set. Let $T{\Lambda^{\prime}}=E^{\prime s}\oplus\langle X\rangle\oplus E^{\prime u}$ be the hyperbolic splitting of $\Lambda^{\prime}$. It is well known that dominated splittings vary continuously and hence $\dim(E^{\prime s})$ and $\dim(E^{\prime u})$ are constant along $\Lambda$, because $\Lambda$ is a chain transitive set. Moreover, up to possibly reducing $U$, we can assume $dim(E^{\prime s})$ and $\dim(E^{\prime u})$ are constant along $\Lambda^{\prime}$ because $d^{H}(\gamma_{n},\Lambda)\to 0$. Also, the continuity of the splitting and the compactness of $\Lambda^{\prime}$ implies that the angles between the fibers $E^{\prime s}$, $\langle X\rangle$ and $E^{\prime u}$ are uniformly limited away from zero.
[…]
for $n$ sufficiently large. Consequently, $\Lambda$ contains a periodic orbit. Then, the result follows from item (2) of Lemma 3.2. ∎
Now, we recall some previous results obtained in [29], in order to get the proof of Theorem 3.1 in the singular case. For this, we follow the exposition given in the aforementioned reference. By Lemma 2.2 any singularity $\sigma\in\Lambda$ is hyperbolic. Let consider the hyperbolic splitting $E^{s}_{\sigma}\oplus E^{u}_{\sigma}$ of $\sigma$. Without loss of generality, we can think of $\sigma$ as the origin in $\mathbb{R}^{n}$, and assume that $E^{s}_{\sigma}$ and $E^{u}_{\sigma}$ are perpendicular (which is possible if one changes the metric). In particular, we will assume that $E^{s}_{\sigma}=\mathbb{R}^{s}$ is the $s$-dimensional subspace of $\mathbb{R}^{n}$ with the last nsn-s coordinates being zero. Here $s=\dim E^{s}_{\sigma}$ is the stable index of $\sigma$. Similarly, $E^{u}_{\sigma}$ is the subspace of $\mathbb{R}^{n}$ where the first $s$ coordinates are zero.
Since the vector field $X$ is $C^{1}$, we can take a neighborhood $V_{\sigma}=B_{r}(\sigma)$ with $r$ small enough, such that the flow in $V_{\sigma}$ can be written as
$X_{t}(x)=e^{At}x+C^{1}\text{ small perturbation},$
where $A$ is a matrix with no eigenvalue on the imaginary axis.
For each $x\in V_{\sigma}$, denote by $x^{s}$ its distance to $E^{s}_{\sigma}$ and $x^{u}$ its distance to $E^{u}_{\sigma}$. Then for every $\alpha>0$ small, we define the $\alpha$-cone on the manifold, denote by $D^{i}_{\alpha}(\sigma)$, $i=s,u$, as follows:
$D^{s}_{\alpha}(\sigma)=\{x\in V_{\sigma}:x^{u}<\alpha x^{s}\},\qquad D^{u}_{\alpha}(\sigma)=\{x\in V_{\sigma}:x^{s}<\alpha x^{u}\}.$
Note that the hyperbolic splitting $T_{\sigma}M=E^{s}_{\sigma}\oplus E^{u}_{\sigma}$ can be extended to $V_{\sigma}$ in a natural way: for each $x\in V_{\sigma}$, put $E^{s}(x)$ as the $s$-dimensional hyperplane that is parallel to $E^{s}_{\sigma}$; the same can be done for $E^{u}(x)$. This allows us to consider the $\alpha$-cones $C_{\alpha}(E^{i})$, $i=s,u$, on the tangent bundle of $V_{\sigma}$.
Let us fix some $\alpha>0$ small enough. One can think of the region $V_{\sigma}\setminus(D^{s}_{\alpha}(\sigma)\cup D^{u}_{\alpha}(\sigma))$ as the place where the flow is “making the turn” from the $E^{s}$-cone to the $E^{u}$-cone. For each $x\in U$, define
$t^{+}(x)=\sup\{t>0:X_{[0,t]}(x)\subset V_{\sigma}\},\qquad t^{-}(x)=\sup\{t>0:X_{[-t,0]}(x)\subset V_{\sigma}\}.$
It is easy to check that the orbit segment $X_{(-t^{-},t^{+})}(x)$ contains $x$ and it is contained in $U$. In this way, we write
(2) $x_{e}=X_{-t_{i}^{-}}(x),\qquad x_{o}=X_{t_{i}^{+}}(x)$
Lemma 3.5 (Lemma 3.2 of [29]).
Let $\sigma$ be a hyperbolic singularity for a $C^{1}$ vector field $X$. Then for every $\alpha>0$ small enough, there exists $T_{\alpha}>0$ such that for every $r>0$ small enough and every $x\in V_{\sigma}=B_{r}(\sigma)$, the set
$T(x):=\{t\in(-t^{-},t^{+}):X_{t}(x)\notin D^{s}_{\alpha}(\sigma)\cup D^{u}_{\alpha}(\sigma)\}$
has length smaller than $T_{\alpha}$.
$N_{x}=\{v\in T_{x}M:\langle v,X(x)\rangle=0\}.$
$\psi^{*}_{t}(v)=\frac{\|X(x)\|}{\|X(X_{t}(x))\|}\psi_{t}(v).$
In [6] it is showed that a partially hyperbolic splitting $T_{\Lambda}M=E\oplus F$ of a compact invariant set $\Lambda$ induces a partially hyperbolic splitting $T_{x}N_{x}=\mathcal{E}_{x}\oplus\mathcal{F}_{x}$ of the normal subspace $N_{x}$ for any $x\in\Lambda\setminus Sing(X)$, where $\mathcal{E}_{x}=\pi(E_{x})$ and $\mathcal{F}_{x}=\pi_{x}(F_{x})$ for every $x\in\Lambda$, where $\pi$ denotes the orthogonal projection. Indeed, since $F_{x}$ contains the flow direction, it follows that $\pi(v)\in F_{x}$ for any $v\neq X(x)$ in $F_{x}$, hence $\mathcal{F}_{x}\subset F_{x}$ because $\pi$ is linear. So, since the splitting $E\oplus F$ is dominated for $\Phi_{t}$, we deduce the dominance of the splitting $\mathcal{E}\oplus\mathcal{F}$. The uniform contracting property for $\pi(E)$ follows because $\|\pi(v)\|\leq\|v\|$ for every $v\in E$. It is important to observe that the above property remains valid for the rescaled linear Poincaré flow $\psi^{*}_{t}$.
Recall the notion of $(\eta,T,\mathcal{F})$-backward contracting points.
Definition 3.6.
Let $X\in X^{1}(M)$, $\Lambda$ a compact invariant set of $X$, and $\mathcal{F}\subset N_{\Lambda\setminus Sing(X)}$ an invariant bundle of the rescaled linear Poincaré flow $\psi^{*}_{t}$. For $0<\eta<1$ and $T_{0}>0$, we say that an orbit segment $X_{[-T,0]}(x)$, $x\in\Lambda\setminus Sing(X)$, is $(\eta,T_{0})$-backward contracting for the bundle $\mathcal{F}$ if there is a partition $0=t_{0}<t_{1}<t_{2}<\cdots<t_{k}=T$, where $t_{i+1}-t_{i}\in[T_{0},2T_{0}]$ for every $i=0,1,\ldots,k-1$ such that
$\prod_{i=0}^{m-1}\|\psi^{*}_{-(t_{i+1}-t_{i})}(X_{t_{i+1}}(x))|_{\mathcal{F}(X_{t_{i+1}}(x))}\|\leq\eta^{m}.$
On the other hand, we say that $x\in\Lambda\setminus Sing(X)$ is a $(\eta,T_{0})$-backward hyperbolic point for the bundle $\mathcal{F}$ if there is a partition $0=t_{0}<t_{1}<t_{2}<\cdots<t_{k}<\cdots$, where $t_{i+1}-t_{i}\in[T_{0},2T_{0}]$ for every $i=0,1,\ldots$ such that
$\prod_{i=0}^{k-1}\|\psi^{*}_{-(t_{i+1}-t_{i})}(X_{t_{i+1}}(x))|_{\mathcal{F}(X_{t_{i+1}}(x))}\|\leq\eta^{k},\quad\forall k\geq 1.$
The next lemma explain the behavior of the orbit segments associated to points in the cones $D_{\alpha}^{*}(\sigma)$, =s,u*=s,u, for every $\sigma\in Sing(X|_{\Lambda})$, with respect to the scaled LPF.
Lemma 3.7 (Lemma 3.9 of [29]).
Let $\sigma$ be a Lorenz-like singularity, then there exists $\eta\in(0,1)$, $T_{0}>0$, $\alpha_{1}>0$ and $r>0$ such that if $\alpha<\alpha_{1}$ and $x$ is a periodic orbit such that the orbit segment $X_{[0,T]}(x)$ is contained in $B_{r}(\sigma)\cap\overline{D_{\alpha}^{s}(\sigma)}$, then the orbit segment is $(\eta,T_{0})$-backward contracting. If the orbit segment is in $\overline{D_{\alpha}^{u}(\sigma)}$, then it does not have any sub-segment that is backward contracting.
Finally, we present the next lemma, which allows us to find a size of the cone field on the $\alpha$-cone on the manifold in order to get transverse intersections between invariant manifolds.
Lemma 3.8 (Lemma 3.11 of[29]).
Let $\sigma$ be a Lorenz-like singularity. Then for $r>0$ small enough, for every $\eta\in(0,1)$, $T_{0}>0$ there exists $\alpha_{2}>0$ such that for all $\alpha<\alpha_{2}$, if $y\in B_{r}(\sigma)\cap Cl(D^{c}_{\alpha}(\sigma))$ is a $(\eta,T_{0})$-backward hyperbolic point on its orbit, then
$W^{u}(y)\cap W^{cs}(\sigma)\neq\varnothing.$
Remark 3.9.
In our case, since $Sing(X)_{\Lambda}$ is finite, the positive nummber
$\alpha=\min_{\sigma\in Sing(X|_{\Lambda})}\{\alpha_{1}(\sigma),\alpha_{2}(\sigma)\})>0$
satisfies the conclusion of Lemma 3.7 and Lemma 3.8 for every $\sigma\in Sing(X|_{\Lambda})$.